OpenStudy (anonymous):

Please Help Ill Give Medal and Fan U!!!!!!! The area of a rectangle piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show steps of our work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (paki):

Area of a rectangle is A=l∗w where w is the width and l is the length. We know the area of this rectangle is 280, so 280=l∗w. Next, we know that if we change the length to be 5 meters less, and the width to be 1 meter more, they will be equal (because the shape will be a square). That implies that l−5=w+1 Using the two equations 280=l∗w l−5=w+1 you can work out the values of the length and width.. Area of a rectangle is A=l∗w where w is the width and l is the length. We know the area of this rectangle is 280, so 280=l∗w. Next, we know that if we change the length to be 5 meters less, and the width to be 1 meter more, they will be equal (because the shape will be a square). That implies that l−5=w+1 Using the two equations 280=l∗w l−5=w+1 you can work out the values of the length and width